# Decision-focused predictions via pessimistic bilevel optimization: complexity and algorithms

### v4

@@ Abstract
Dealing with uncertainty in optimization parameters is an important and longstanding challenge. Typically, uncertain parameters are predicted accurately, and then a deterministic optimization problem is solved. However, the decisions produced by this so-called predict-then-optimize procedure can be highly sensitive to uncertain parameters. In this work, we contribute to recent efforts in producing decision-focused predictions, i.e., to build predictive models that are constructed with the goal of minimizing a regret measure on the decisions taken with them. We begin by formulating the exact expected regret minimization as a pessimistic bilevel optimization model. Then, we establish NP-completeness of this problem in a heavily restricted case. Using duality arguments, we reformulate it as a non-convex quadratic optimization problem. Finally, leveraging the quadratic reformulation, we show various computational techniques to achieve empirical tractability. We report extensive computational results on shortest-path and bipartite matching instances with uncertain cost vectors. Our results indicate that our approach can improve training performance over the approach of Elmachtoub and Grigas (2022), a state-of-the-art method for decision-focused learning.
Keywords:
predict and optimize pessimistic bilevel optimization non-convex quadratics
@@ 1 Introduction
In recent years, decision-focused learning (DFL) (also known as smart predict-then-optimize (SPO)) approaches, in which prediction and optimization tasks are integrated into the learning process, have received significant attention. In this approach, a machine learning model is specifically trained to enhance the effectiveness of the whole decision-making process. This involves combining, during the training phase, the prediction and the optimization in one single model.
@@
As we will see in the following sections, finding the model $m(\omega,x)$ that minimizes the regret can be formulated as a pessimistic bilevel optimization problem. In fact, authors in Elmachtoub and Grigas, (2022) define the unambiguous-SPO loss function as a pessimistic regret: among all the solutions that minimize predicted objective function, it penalizes the one that hurts the most when it is evaluated with the true cost vector.
Unfortunately, the complexity of finding these models and scalability are two major roadblocks to this DFL approach. Consequently, the literature has mainly focused on stochastic gradient-based approaches via approximations of the loss function through a surrogate convex loss function and/or solving a relaxed optimization problem (see Mandi et al., (2024)). Here, we follow a different path, take a step back, and focus on carefully studying the mathematical object behind the exact expected regret minimization. Our main hypothesis is that by understanding the mathematical object and designing new methods for solving the pessimistic bilevel optimization problem, better predictions and better algorithms can be developed.
The contributions of this work are the following: i. we formulate the expected regret minimization problem as a pessimistic bilevel optimization problem; ii. we prove that the problem is NP-complete in a heavily restricted setting, which improves on the known hardness results; iii. we reformulate the bilevel pessimistic formulation as a non-convex quadratically-constrained quadratic program (QCQP), which can be tackled by current optimization technology for moderate sizes; iv. we propose heuristics to improve the solution procedure based on the quadratic reformulation; 5) we conduct a comprehensive computational study on shortest path and bipartite matching instances. An early version of this work was published as a short paper in the conference proceedings of CPAIOR 2024 (Bucarey et al.,, 2024).
@@ 1.1 Problem setting
$\mbox{Regret}(\mathcal{D},\omega):=\qquad\qquad\qquad\qquad$
$\max_{v}\quad$ $\frac{1}{N}\sum_{i\in[N]}\left(c^{i}{}^{\top}v^{i}-z^{*}(c^{i})\right),$ (2)
@@
Here, we are comparing the true optimal value $z^{*}(c^{i})$ with the value $c^{i}{}^{\top}v^{i}$, which is the “true cost” of a solution that is optimal for the prediction $m(\omega,x^{i})$. To find the values of $\omega$ that minimize the regret (1.1), we must solve the following pessimistic bilevel optimization problem.
@@ 1.2 Importance of the pessimistic approach
To illustrate why the pessimistic approach is relevant (as opposed to taking an optimistic approach), we consider the following example of a linear optimization problem with two variables:
$P(c):\quad\min_{v}\quad$ $c_{1}v_{1}+c_{2}v_{2}$
s.t. $v_{1}+v_{2}\leq 1$
$v_{1},v_{2}\geq 0.$
Suppose that we have the two following observations of the true cost vectors with one single feature:
$(x^{1},c^{1})=$ $\,\left(\begin{pmatrix}1\\ 0\end{pmatrix},\begin{pmatrix}-4\\ -3.5\end{pmatrix}\right),$
$(x^{2},c^{2})=$ $\,\left(\begin{pmatrix}0\\ -2\end{pmatrix},\begin{pmatrix}-2\\ -3\end{pmatrix}\right),$
where each feature vector corresponds to the attribute regarding one coordinate. The goal is to estimate a simple regression model $m(\omega,x_{j})=\hat{c}_{j}(\omega)=\omega_{0}+\omega_{1}x_{j}$, one per coordinate. Note that we can assume, without loss of generality, $|\omega_{i}|\leq 1$ for $i=0,1$.
The optimistic version is obtained dropping the “$\max_{v\in V^{*}(\hat{c}^{i}(\omega))}$” in (4); this effect can be equivalently achieved by setting $\omega_{0}=\omega_{1}=0$, since $V^{*}((0,0))=V$ (i.e., every $v\in V$ is optimal). The optimistic version then “guesses” the right optimal solution, and thus, the optimistic regret equals 0. However, the values for this solution evaluated in the pessimistic version are 4 for cost vector $c^{1}$ and 3 for $c^{2}$, leading to a regret of $3.5$. Thus, this “optimistic regret” does not adequately measure the sensitivity of the decision with respect to the predictive model.
Note that the classical linear regression performs better than the optimistic approach in terms of pessimistic regret. Actually, the minimum squared error estimator of this regression is $\omega_{0}=-3.156$, $\omega_{1}=0.253$, producing a regret of 0.5.
By solving the bilevel pessimistic formulation exactly, we obtain a regret of 0.25. An optimal solution here is $\omega_{0}=-1$, $\omega_{1}=0.125$: with these values, the predicted costs are $\hat{c}^{1}=(-0.875,-1)^{\top}$, $\hat{c}^{2}=(-1,-1.125)^{\top}$. Note that this is not the only optimal solution for this example. If we consider $\omega_{0}=-3.125$, $\omega_{1}=0.1$ also returns a pessimistic regret equal to 0.25. We depict the different solutions and the observations in Figure 1, where the green stars correspond to each coordinate of observation 1 ($x^{1},c^{1}$) and black squares correspond to observation 2 ($x^{2},c^{2}$).
Figure 1: Linear regression and regret minimizers in the numerical example of Section 1.2. Black squares and green stars correspond to the observations. Dotted lines correspond to the different solutions for the linear model $m(\omega,x_{j})$.
1.3 Related work
@@
Bilevel optimization problems are hierarchical ones: they consist of an optimization problem (called upper-level or leader problem) that contains in its constraints optimality conditions of other problems, called lower-level or follower problems (Dempe and Zemkoho, (2020)). In the presence of several optimal solutions at the lower-level problems, the way that the solution is chosen opens two approaches: the optimistic approach, also known as the strong solution, in which the solution chosen is the one that favors the upper-level optimization problem; and the pessimistic, also known as the weak solution, that chooses the one that worsens the most the upper-level objective.
Optimistic bilevel optimization is NP-hard even when the upper and lower-level problems are linear. Even having compact/efficient formulations can be a difficult task (Kleinert et al., (2020)). See Kleinert et al., (2021) for a survey on mixed-integer programming techniques in bilevel optimization. The necessity of establishing the pessimistic solution for bilevel optimization problems was first raised by Leitmann, (1978) and studied in several articles (e.g., Loridan and Morgan, (1996); Aboussoror and Mansouri, (2005)). It is a common belief that the pessimistic approach is much more difficult than the optimistic one (Zeng, (2020)). We refer to readers to Wiesemann et al., (2013); Liu et al., (2018, 2020) for comprehensive surveys of theoretical background and methods for bilevel pessimistic optimization problems. We note that the hardness results in bilevel optimization rule out the existence of efficient general-purpose techniques that could be applied in our setting.
@@
As mentioned earlier, and to the best of our knowledge, the existing methods developed for training decision-focused predictions are primarily based on estimating the gradient of how predictions impact a specific loss function. The main challenge in these approaches is to estimate the changes in the optimal solution with respect to the model parameters, known as optimization mapping. According to the recent survey by Mandi et al., (2024), these methods can be classified into four categories: i. those that compute the gradient of the optimization mappings analytically, as in Amos and Kolter, (2017); or ii. solving a smooth version of the optimization mapping, as seen in Wilder et al., (2019); Mandi and Guns, (2020); Ferber et al., (2020); iii. ; or smoothing the optimization mapping by random perturbations Pogancic et al., (2020); Niepert et al., (2021) ; or iv. those solving a surrogate loss function that approximates regret, as discussed in Elmachtoub and Grigas, (2022); Mulamba et al., (2021); Mandi et al., (2022).
Our study diverges from these approaches. Instead, to find regret-minimizing models, we leverage the pessimistic bilevel optimization formulation (4), and by employing duality arguments, we formulate it as a single-level non-convex quadratic model. A related method for minimizing the expected regret was proposed by Jeong et al., (2022): they cast the problem, not explicitly, as an optimistic bilevel optimization problem and use a symbolic reduction to solve this problem. However, as mentioned in Elmachtoub and Grigas, (2022) and discussed in Section 1.2, casting this problem as an optimistic bilevel optimization problem may lead to undesirable predictions.
In terms of the complexity of (4), the results of Elmachtoub and Grigas, (2022) imply that this problem is NP-hard. They show that minimizing the regret function generalizes the minimization of the 0-1 loss function, which is known to be NP-hard (Ben-David et al.,, 2003). Here, we improve on this hardness result by showing that in a very restrictive case, the problem remains NP-hard. We also complement this by showing membership in NP of a decision version of (4).
@@ 2 Complexity results
In this section, we focus on studying the complexity of (4) in a basic setting: when the predictive model $m(\omega,x)$ is linear. Technically, $x$ are feature vectors, but slightly abusing notation, we will treat them as matrices in this section. This allows to write $m(\omega,x)=x\omega$ for a linear predictive model.
As mentioned in the previous section, the results of Elmachtoub and Grigas, (2022) imply that (4) is NP-hard since it generalizes the minimization of the 0-1 loss function, which is known to be NP-hard (Ben-David et al.,, 2003).
In this section, we provide a sharper hardness result by showing that (4) is NP-hard even if we restrict to a linear nominal problem with a bounded feasible region and only one observation. We also show membership in NP, thus establishing NP-completeness of the problem.
Remark 1
The reduction in Elmachtoub and Grigas, (2022) of the 0-1 loss minimization to (4) does not provide NP-hardness in the case we consider here since the former can be solved in polynomial time when there is only one observation (Blumer et al.,, 1989). This provides an interesting separation (complexity-wise) of problem (4) and the minimization of the 0-1 loss.
Remark 2
While many NP-hardness results are known in bilevel optimization (e.g. Audet et al., (1997); Hansen et al., (1992); Vicente et al., (1994); Buchheim, (2023)), they are not directly applicable to (4) due to the following reasons:
  • Problem (4) is a pessimistic bilevel optimization problem.
  • The leader variables $\omega$ are unrestricted.
  • The leader and follower variables interact non-linearly in the follower’s problem, even when $m(\omega,x)$ is a linear function.
Adapting the existing bilevel optimization reductions to our setting seems like a highly non-trivial task, which is why we opted for a new custom proof.
Let us consider the following decision problem, associated with a simplified version of the optimization problem (4).
Definition 1
@@
$\max_{v^{i}\in V^{*}(x^{i}\omega)}\quad\frac{1}{N}\sum_{i\in[N]}c^{i}{}^{\top}v^{i}\quad\leq\quad M$ (6)
We note that (6) is obtained from (4) by restricting $m(\omega,x)=x\omega$, and removing the constant part of the objective function $z^{*}(c^{i})$; ; these terms can be computed efficiently, as they are only linear programs..
Theorem 2.1
@@
Proof
[…]
Fix an instance of 3SAT and take $\{C_{j}\}_{j=1}^{m}$ the set of clauses and $u_{k}$, $k\in[n]$, the set of variables. For each clause $j$ we denote $j_{1},j_{2},j_{3}$ the indices of the 3 variables appearing in $C_{j}$. We also define $f_{j}(u_{j_{\ell}})$ as
$f_{j}(u_{j_{\ell}})=\left\{\begin{array}[]{ll}u_{j_{\ell}}&\text{if $C_{j}$ contains $u_{j_{\ell}}$ as a literal}\\ 1-u_{j_{\ell}}&\text{if $C_{j}$ contains $\overline{u_{j_{\ell}}}$ as a literal}\end{array}\right.$
Using this, we let $V\subseteq\mathbb{R}^{n+m}$ be the polytope defined by
$y_{j}$ $\geq f_{j}(u_{j_{\ell}})\quad\forall j\in[m],\,\ell=1,2,3$
$y$ $\in[0,1]^{m}$
$u$ $\in[0,1]^{n}$
[…]
$c^{1}=(\underbrace{0,\ldots,0}_{n},\underbrace{-1,\ldots,-1}_{m})\in\mathbb{R}^{n+m}.$
Additionally, we define $x^{1}$ as the following $(n+m)\times(n+m)$ matrix
$x^{1}=\begin{bmatrix}I_{n}&0\\ 0&0\end{bmatrix}$
where $I_{n}$ is the $n\times n$ identity matrix and the zeros are stand-ins for all-zero matrices of appropriate dimensions. This way, $x^{1}\omega=(\omega_{u},0)$ where $\omega_{u}$ is the sub-vector of $\omega$ consisting of its first $n$ entries. Finally, we set $M=-m$.
[…]
$\max\quad$ $-\sum_{j=1}^{m}y_{j}$ (7a)
s.t $(u,y)\in\arg\min\,\big{\{}\omega_{u}^{\top}u\,:\,$ (7b)
$\qquad y_{j}\geq f_{j}(u_{j_{\ell}})\quad\forall j\in[m],\,\ell=1,2,3,$ (7c)
$\qquad y\in[0,1]^{m},$ (7d)
$\qquad u\in[0,1]^{n}\big{\}}$ (7e)
Clearly, (7) can be constructed in polynomial time. We now show that 3SAT outputs “Yes” if and only there exists $\omega$ such that the value of (7) is at most $-m$.
If the 3SAT formula is satisfiable, we can take $\hat{u}$ an assignment of the $u$ variables that makes all clauses true. With this, we define
$(\omega_{u})_{k}=\left\{\begin{array}[]{ll}-1&\text{if }\hat{u}_{k}=1\\ 1&\text{if }\hat{u}_{k}=0\end{array}\right.$
and we observe that $(u,y)$ satisfies (7b) if and only if $u=\hat{u}$ and $y$ satisfies that $y_{j}=1$ when clause $C_{j}$ is satisfied by $\hat{u}$, or $y_{j}\in[0,1]$ otherwise. Since $\hat{u}$ was a satisfiable assignment, this means that all $m$ entries of $y$ will be 1, and thus
$-\sum_{j=1}^{m}y_{j}\leq-m.$

### v5

@@ Abstract
Dealing with uncertainty in optimization parameters is an important and longstanding challenge. Typically, uncertain parameters are predicted accurately, and then a deterministic optimization problem is solved. However, the decisions produced by this so-called predict-then-optimize procedure can be highly sensitive to uncertain parameters. In this work, we contribute to recent efforts in producing decision-focused predictions, i.e., to build predictive models that are constructed with the goal of minimizing a regret measure on the decisions taken with them. We begin by formulating the exact expected regret minimization as a pessimistic bilevel optimization model. Then, we show computational complexity results of this problem, including its membership in NP. In combination with a known NP-hardness result, this establishes NP-completeness and discards its hardness in higher complexity classes. Using duality arguments, we reformulate it as a non-convex quadratic optimization problem. Finally, leveraging the quadratic reformulation, we show various computational techniques to achieve empirical tractability. We report extensive computational results on shortest-path and bipartite matching instances with uncertain cost vectors. Our results indicate that our approach can improve training performance over the approach of Elmachtoub and Grigas (2022), a state-of-the-art method for decision-focused learning.
Keywords: Predict-and-optimize , Pessimistic bilevel optimization, Non-convex quadratics
@@ 1 Introduction
In recent years, decision-focused learning (DFL) (also known as smart predict-then-optimize (SPO)) approaches, in which prediction and optimization tasks are integrated into the learning process, have received significant attention. In this approach, a machine learning model is specifically trained to enhance the effectiveness of the whole decision-making process. This involves combining, during the training phase, the prediction and the optimization in a single model.
@@
As we will see in the following sections, finding the model $m(\omega,x)$ that minimizes the regret can be formulated as a pessimistic bilevel optimization problem. In fact, authors in elmachtoub2022smart define the unambiguous-SPO loss function as a pessimistic regret: among all the solutions that minimize the predicted objective function, it penalizes the one that hurts the most when it is evaluated with the true cost vector.
Unfortunately, the complexity of finding these models and scalability are two major roadblocks to this DFL approach. Consequently, the literature has mainly focused on stochastic gradient-based approaches via approximations of the loss function through a surrogate convex loss function and/or solving a relaxed optimization problem (see mandi2023decision ). Here, we follow a different path, take a step back, and focus on carefully studying the mathematical object behind the exact expected regret minimization. Our main hypothesis is that by understanding the mathematical object and designing new methods for solving the pessimistic bilevel optimization problem, better predictions and better algorithms can be developed.
The contributions of this work are the following: (i) we formulate the expected regret minimization problem as a pessimistic bilevel optimization problem; (ii) we prove that the problem belongs to the NP complexity class, which settles NP-completeness and makes it unlikely for it to be higher in the polynomial hierarchy (i.e., the problem is not $\Sigma_{2}^{P}$-hard, unless the hierarchy collapses); (iii) we show that, under mild assumptions, determining if the regret is 0 is polynomial-time solvable; (iv) we reformulate the bilevel pessimistic formulation as a non-convex quadratically-constrained quadratic program (QCQP), which can be tackled by current optimization technology for moderate sizes; (v) we propose heuristics to improve the solution procedure based on the quadratic reformulation; (vi) we conduct a comprehensive computational study on shortest path and bipartite matching instances. An early version of this work was published as a short paper in the conference proceedings of CPAIOR 2024 (bucarey2024decision, ), where the single-level reformulation and a local-search heuristic were presented. We extend this work by providing theoretical results regarding the complexity of this problem and an alternate descent direction method. We finally extend the experiments by including the bipartite matching problem in the computational study.
We note that, in parallel to the preparation of this extended version, the work of jimenez2025pessimistic also proposed to use a pessimistic bilevel optimization approach for DFL. The method proposed is considerably different, as jimenez2025pessimistic proposes a column-and-constraint generation and a branch-and-cut approach. A close comparison is subject of future work.
@@ 1.1 Problem setting
$\mbox{Regret}(\mathcal{D},\omega):=\,\max_{v}\quad$ $\frac{1}{N}\sum_{i\in[N]}\left(c^{i}{}^{\top}v^{i}-z^{*}(c^{i})\right),$ (2)
@@
Here, we are comparing the true optimal value $z^{*}(c^{i})$ with the value $c^{i}{}^{\top}v^{i}$, which is the “true cost” of a solution that is optimal for the prediction $m(\omega,x^{i})$. To find the values of $\omega$ that minimize the regret (2), we must solve the following pessimistic bilevel optimization problem.
@@ 1.2 Importance of the pessimistic approach
To illustrate the relevance of the pessimistic approach (in contrast to the optimistic one), we consider the following linear optimization problem with two variables:
$P(c):\quad\min_{v}\quad c_{1}v_{1}+c_{2}v_{2}\quad\text{s.t.}\quad v_{1}+v_{2}\leq 1,\quad v_{1},v_{2}\geq 0.$
Suppose we observe the following three instances of the true cost vectors, each associated with a single feature value:
$(x^{1},c^{1})=\left(0,\begin{pmatrix}-3,-2\end{pmatrix}\right)^{\top},\,(x^{2},c^{2})=\left(1,\begin{pmatrix}-2,-5\end{pmatrix}\right),\,(x^{3},c^{3})=\left(2,\begin{pmatrix}-2,0\end{pmatrix}\right)^{\top}$.
Our goal is to estimate a linear regression model of the form $m(\omega,x)_{j}=\hat{c}_{j}(\omega)=\omega_{0j}+\omega_{1j}x,\quad\text{for }j=1,2.$
We begin by analyzing the optimistic version, which corresponds to dropping the outer maximization $\max_{v\in V^{*}(\hat{c}^{i}(\omega))}$ in (4). This is equivalent to setting $\omega_{0j}=\omega_{1j}=0$ for both $j=1,2$, since in that case $\hat{c}=(0,0)$ and the solution set becomes $V^{*}((0,0))=V$ (i.e., every feasible $v\in V$ is optimal).
Thus, the optimistic model “guesses” the correct decision in all cases, resulting in zero regret. However, the pessimistic model evaluates the regret over all possible solutions in $V^{*}((0,0))$ and selects the worst-case one—in this case, $v_{1}=v_{2}=0$. The regrets under this decision are 3, 5, and 2 for observations 1, 2, and 3, respectively, yielding an average regret of $\frac{10}{3}$. Thus, this “optimistic regret” does not adequately measure the sensitivity of the decision with respect to the predictive model.
In contrast, a classical linear regression performs better under the pessimistic regret criterion. The minimum squared error solution for this regression is
$\omega=\begin{pmatrix}-2.83&0.50\\ -3.33&0.99\end{pmatrix}$,
which induces the correct decision for observation 3 and results in an average regret of $\frac{4}{3}$.
Solving the bilevel pessimistic formulation exactly leads to an even lower regret of $\frac{1}{3}$, with an optimal regression model
$\omega=\begin{pmatrix}-1&-1\\ -4&1\end{pmatrix}$.
This solution correctly predicts the optimal decision for observations 2 and 3 (those with the highest regret) and only fails on observation 1.
We also evaluate the SPO+ estimator proposed by elmachtoub2022smart , computed via a linear programming formulation. This method yields an average regret of 1.
Figure 1 summarizes the solutions and outcomes for each approach.
Figure 1: Linear regression, SPO+ estimator from elmachtoub2022smart and an exact regret minimizer for the numerical example of Section 1.2. Red circles and squares represent the input data. Dash-dotted lines correspond to the different solutions for the linear model $m(\omega,x)$.
1.3 Literature Review
@@
Bilevel optimization problems are hierarchical ones: they consist of an optimization problem (called upper-level or leader problem) that contains in its constraints optimality conditions of other problems, called lower-level or follower problems dempe2020bilevel . In the presence of several optimal solutions at the lower-level problems, the way that the solution is chosen opens two approaches: the optimistic approach, also known as the strong solution, in which the solution chosen is the one that favors the upper-level optimization problem; and the pessimistic, also known as the weak solution, that chooses the one that worsens the most the upper-level objective.
Optimistic bilevel optimization is NP-hard even when the upper and lower-level problems are linear. Even having compact/efficient formulations can be a difficult task kleinert2020there . See kleinert2021survey for a survey on mixed-integer programming techniques in bilevel optimization. The necessity of establishing the pessimistic solution for bilevel optimization problems was first raised by leitmann1978generalized and studied in several articles (e.g., loridan1996weak ; aboussoror2005weak ). Until recently, it was a common belief that the pessimistic approach is much more difficult than the optimistic one. However, it has been established that a pessimistic linear problem can be transformed equivalently into an optimistic one. henke2025coupling ; zeng2020practical We refer to readers to wiesemann2013pessimistic ; liu2018pessimistic ; liu2020methods for comprehensive surveys of theoretical background and methods for bilevel pessimistic optimization problems. We note that the hardness results in bilevel optimization rule out the existence of efficient general-purpose techniques that could be applied in our setting. In addition, the equivalence between pessimistic and optimistic was established for the linear case; as we will see, our model is non-linear.
@@
As mentioned earlier, and to the best of our knowledge, the existing methods developed for training decision-focused predictions are primarily based on estimating the gradient of how predictions impact a specific loss function. The main challenge in these approaches is to estimate the changes in the optimal solution with respect to the model parameters, known as optimization mapping. According to the recent survey by mandi2023decision , these methods can be classified into four categories: i. those that compute the gradient of the optimization mappings analytically, as in amos2017optnet ; or ii. solving a smooth version of the optimization mapping, as seen in wilder2019melding ; mandi2020interior ; ferber2020mipaal ; or iii. smoothing the optimization mapping by random perturbations PogancicPMMR20 ; niepert2021implicit ; or iv. those solving a surrogate loss function that approximates regret, as discussed in elmachtoub2022smart ; mulamba2021contrastive ; mandi2022decision .
Our study diverges from these approaches. Instead, to find regret-minimizing models, we leverage the pessimistic bilevel optimization formulation (4), and by employing duality arguments, we formulate it as a single-level non-convex quadratic model. A related method for minimizing the expected regret was proposed by jeong2022exact : they cast the problem, not explicitly, as an optimistic bilevel optimization problem and use a symbolic reduction to solve this problem. However, as mentioned in elmachtoub2022smart and discussed in Section 1.2, casting this problem as an optimistic bilevel optimization problem may lead to undesirable predictions. The recent work of jimenez2025pessimistic also proposes an exact pessimistic bilevel optimization approach.
In terms of the complexity of (4), the results of elmachtoub2022smart imply that this problem is NP-hard. They show that minimizing the regret function generalizes the minimization of the 0-1 loss function, which is known to be NP-hard (ben2003difficulty, ). Here, we complement this by showing membership in NP of a decision version of (4).
@@ 2 Complexity results
In this section, we focus on studying the complexity of (4) in a basic setting: when the predictive model $m(\omega,x)$ is linear. Since we are predicting cost vectors, we can assume our set of parameters $\omega$ are in matrix form, i.e., $m(\omega,x)=\omega x$.
As mentioned in the previous section, the results of elmachtoub2022smart imply that (4) is NP-hard since it generalizes the minimization of the 0-1 loss function, which is known to be NP-hard (ben2003difficulty, ). In this section, we begin by showing membership in NP, thus establishing NP-completeness of the problem. Note that Buchheim buchheim2023bilevel recently showed that linear bilevel optimization, both optimistic and pessimistic, belongs to NP. However, this result is not directly applicable to (4) since the leader and follower variables interact non-linearly in the follower’s problem, even when $m(\omega,x)$ is a linear function.
2.1 Membership in NP
Let us consider the following decision version of (4).
Definition 1.
The decision problem SIMPLE-REGRET is defined as follows. Given $(c^{i},x^{i})_{i=1}^{N}$ a collection of $N$ rational vectors and matrices, a polytope $V$, and a rational number $M$ decide if there exists $\omega$ such that
$\max_{v^{i}\in V^{*}(\omega x^{i})}\quad\frac{1}{N}\sum_{i\in[N]}(c^{i}{}^{\top}v^{i}-z^{*}(c^{i}))\quad\leq\quad M.$ (5)
We note that (5) is obtained from (4) by restricting $m(\omega,x)=\omega x$.
Theorem 2.
$\mbox{SIMPLE-REGRET}\in\mbox{NP}$
Proof.
It suffices to show that, for any “Yes” instance of SIMPLE-REGRET, there is a polynomially-sized $\omega$ such that (5) holds. Indeed, if we have such an $\omega$, we can simply compute the optimal faces of each of the $N$ lower-level problems (which can be done in polynomial time), and then solve the resulting maximization problem. Let us show that such an $\omega$ exists.
Consider an arbitrary instance of SIMPLE-REGRET, with $V=\{v\,:\,Av\geq b\}$, and suppose there exists $\hat{\omega}$ such that the value of
$\max\quad$ $\frac{1}{N}\sum_{i=1}^{N}(c^{i}{}^{\top}v^{i}-z^{*}(c^{i}))\quad\text{s.t}\quad v^{i}\in\arg\min\,\big{\{}(\hat{\omega}x^{i})^{\top}v\,:\,Av\geq b\big{\}}$
$\max\quad$ $(\rho^{i})^{\top}b\quad\text{s.t}\quad(\rho^{i})^{\top}A=\hat{\omega}x^{i},\quad\rho^{i}\geq 0.$
For each $i$, let us consider $\hat{\rho}^{i}$ in the relative interior of the optimal face of the dual. This means that the optimal face of the $i$-th primal can be described as
$F^{i}(\hat{\omega}):=\{v\in V\,:\,a_{j}^{\top}v=b_{j},\,j\,:\,\hat{\rho}^{i}_{j}>0\}.$
@@ 2.1 Membership in NP
$(\tilde{\rho}^{i})^{\top}A$ $=\tilde{\omega}x^{i}$ $\forall i\in[N]$ (6a)
$\tilde{\rho}^{i}_{j}$ $\geq 1$ $\hat{\rho}^{i}_{j}>0,\,\forall i\in[N]$ (6b)
$\tilde{\rho}^{i}_{j}$ $=0$ $\hat{\rho}^{i}_{j}=0,\,\forall i\in[N]$ (6c)
@@
$\arg\min\{(\tilde{\omega}x^{i})^{\top}v\,:\,Av\geq b\}=$ $\arg\min\{(\tilde{\rho}^{i})^{\top}Av\,:\,Av\geq b\}$
$=$ $\arg\min\left\{\sum_{j:\hat{\rho}^{i}_{j}>0}\tilde{\rho}^{i}_{j}a_{j}^{\top}v\,:\,Av\geq b\right\}$